7. Let the cost function of a monopolist be c()-100+20q. Find the market equilibrium, price elasticity of demand at the equilibrjum, profit, and DWL if (a) market demand is p-100-q (b) market demand is q-120-2p St. 8. There are two identical firms in the market. Each firm has cost function c(g)-24q and the market demand is p-60-q. The two firms are considering whether to produce at cooperative output level (cartel) or to set output level non-cooperatively (Cournot). (a) Find each firms profit if (i) both firms choose to co-operate; 42. Ib. both firms choose not to co-operate; P3b. I ii) one firm chooses to co-operate but another firm chooses not to co-operate. (b) Based on (a), find the market equilibrium if the firms set output le ultaneously?乡136 ls 9. Customers are uniformly located along a street with unit length ( the left end is 0 and the right end is 1). Each of them wants to buy one unit of a good. The transportation cost of customers is c dollars per unit length. Firms have same production cost and the market price is fixed at some level above the production cost. 8
Answer #1

8. A.

i. When they cooperate then the cartel will be as good as monopoly. Thus, they will be producing up to MR = MC

P = 60 - Q

TR = 60Q - Q2

Differentiating TR wrt Q we get

MR = 60 - 2Q

Similarly differentiating TC wrt Q we get

TC = 24Q

MC = 24

Now equate MR = MC

60 - 2Q = 24

2Q = 60 - 24

2Q = 36

Q* = 18 units

P* = $ 42 per unit

Each firms output = 9 units

Profit of each firm = 42*9 -(24*9) = 378 - 216 = $ 162

ii. When they chose not to cooperate

P = 60 - Q1 - Q2

TR1 = 60Q1 - Q12 - Q1Q2

MR1 = 60 - 2Q1 - Q2

MC = 24

Therefore, 60 - 2Q1 - Q2 = 24

2Q1 - Q2 36

2Q1- 36 - Q2

36- Q2

Similarly, solve for Q2

TR2 = 60Q2 -Q1 Q2 - Q22

MR2 = 60 - Q1 - 2Q2

60 - Q1 - 2Q2-24

Plug in Q1

60 _ (18-0.5Q2)-2Q2-24

60-18 + 0.502-292 24

42-1.524

large 1.5Q_2 = 42 - 24

1.592 = 18

5 1 2

Q) 12

Similarly, Q1 = 18 - 0.5*12 = 12 units

P = 60 - 12 - 12 = $ 36 per unit

Profit of both firms = 12*36 - 24*12 = $ 144

iii. When one firm cooperates and other don't

Let us assume firm 1 abides by the agreement thus as per agreement the firm will produce only 9 units(Refer part i).

Firm 2 cheat then the reaction function will be

36-Q1

Plug in Q1 and solve for Q2

large Q_2 = rac{36 - 9}{2} = 13.5

Total output,Q = Q1 +Q2 = 9 + 13.5 = 22.5

P = 60 - Q

P = 60 - 22.5

P = $ 37.50 per unit

Profit of firm 1= 37.50*9 - 24*9 = $ 121.50

Profit of Firm 2 = 37.50*13.5 - 24*13.5 = $ 182.25

B. When the firms set the output level simultaneously then the output will be 12 each and price will be $36

Please contact if having any query thank you.